xdy-ydx=x(x^2-y^2)^1/2


1
Expert's answer
2021-05-07T09:43:07-0400

"Given\ differential\ equation \ is \ xdy-ydx=x(\sqrt{x^2-y^2})...(1)\\
Dividing \ the \ equation \ in \ (1) \ with \ \sqrt{x^2-y^2} \, we \ get \\
\frac{x}{\sqrt{x^2-y^2}}\frac{dy}{dx}-\frac{y}{\sqrt{x^2-y^2}}=x\\
Substituting \ y=vx, and \ \frac{dy}{dx}=v+x\frac{dv}{dx}, we \ get \\
\frac{x}{\sqrt{x^2-v^2x^2}}(v+x\frac{dv}{dx})-\frac{vx}{\sqrt{x^2-v^2x^2}}=x\\
\Rightarrow \frac{x}{\sqrt{x^2(1-v^2)}}(v+x\frac{dv}{dx})-\frac{vx}{\sqrt{x^2(1-v^2)}}=x\\
\Rightarrow \frac{x}{x\sqrt{(1-v^2)}}(v+x\frac{dv}{dx})-\frac{vx}{x\sqrt{(1-v^2)}}=x\\
\Rightarrow \frac{1}{\sqrt{(1-v^2)}}(v+x\frac{dv}{dx})-\frac{v}{\sqrt{(1-v^2)}}=x\\
(by \ cancellation)\\
\Rightarrow \frac{v}{\sqrt{(1-v^2)}}+(\frac{dv}{dx})\frac{x}{\sqrt{(1-v^2)}}-\frac{v}{\sqrt{(1-v^2)}}=x\\
\Rightarrow (\frac{dv}{dx})\frac{x}{\sqrt{(1-v^2)}}=x (by \ cancellation \ law)\\
\Rightarrow (\frac{dv}{dx})\frac{1}{\sqrt{(1-v^2)}}=1 (by \ cancellation \ law)\\
Separating \ the \ variables \ v \ and \ x\\
we \ get \ \frac{dv}{\sqrt{(1-v^2)}}=1 dx\\
Integrating \ on \ both \ sides, \ we \ get \\
\int\frac{dv}{\sqrt{(1-v^2)}}=\int 1 dx\\
\Rightarrow sin^{-1}(v)=x+C, where \ C \ is \ a \ constant \ of\ integration\\
Substituting \ back, \ v= \frac{y}{x}, we \ get \\
sin^{-1}(\frac{y}{x})=x+C\\
Taking \ sin \ on \ both \ sides \ we \ get \\
sin(sin^{-1}(\frac{y}{x}))=sin (x+C)\\
\therefore \frac{y}{x}=sin (x+C)"